# Frontiers of Generative AI for Network Optimization: Theories, Limits, and Visions

### v1

@@ V-A1 Bound of Solution Generation
Our goal is to relate the KL divergence between the GDM-generated solution distribution and the true target solution distribution to the expected objective value of the generated solutions. Due to the asymmetry of KL divergence, we adopt the forward KL divergence, rather than the reverse KL used in Eq. 3, to properly express the expected objective value under the generated distribution.
Specifically, we construct a target solution distribution induced by the objective function, denoted as $p_{0}(\mathbf{y}|\mathbf{x})\propto e^{-\alpha f(\mathbf{y},\mathbf{x})}$. Specifically, let $p_{0}(\mathbf{y}|\mathbf{x})=\frac{e^{-\alpha f(\mathbf{y},\mathbf{x})}}{Z(\mathbf{x})}$, where $Z(\mathbf{x})$ serves as the normalization constant or partition function, and $\alpha>0$ is a temperature parameter that controls the preference for optimal solutions. For large $\alpha$, the distribution concentrates around the optimal solutions; as $\alpha$ approaches zero, the distribution becomes nearly uniform.
[…]
@@
[…]
$\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})}$ $\left[f(\mathbf{y},\mathbf{x})\right]=$ (5)
Based on Eq. 5, we can conclude that
@@
In Eq. 6, $\mathbf{y}^{*}$ denotes the optimal solution corresponding to the given $\mathbf{x}$. The normalization constant $Z$ depends only on the target solution distribution and $\mathbf{y}$, and is thus fixed under the given $\mathbf{x}$. The entropy term captures the uniformity of the model’s output distribution over solutions, where higher entropy implies a more uniform distribution, indicating greater uncertainty and lower expected quality (i.e., higher objective value) of the generated solution. The parameter $\alpha$ characterizes both the distance between modes and the number of modes in the high-quality region of the target distribution. Intuitively, it reflects the complexity of the optimization problem. When $\alpha$ is small, the target distribution is more complex (i.e., more modes or closely spaced modes), implying higher problem complexity and a looser bound on the right side of Eq. 6. The KL divergence term is always non-negative. Although Eq. 6 is derived using the forward KL divergence, which differs from Eq. 3, the latter can still be interpreted as providing an upper bound for this term.
In summary, Eq. 6 highlights several key factors that influence the generalization ability of conditional GDMs in one-shot optimization solution generation. These include the complexity of the optimization problem (reflected in the complexity of the target solution distribution), the convergence quality of the model (reflected in the entropy of the predicted distribution), the prediction accuracy (reflected in the KL divergence), and the factors affecting this KL divergence—namely, the number of training samples, model size, and training duration.
@@ V-A2 Bound of Policy Decision Generation
DRL trains a policy that takes actions based on the input state to receive rewards, aiming to maximize the expected cumulative discounted return. Our objective is to leverage the KL divergence bound between the optimal policy and the learned (predicted) policy, be described by Eq. 3, to quantify the gap between their corresponding value functions.
Consider an infinite horizon MDP =$\mathcal{M}=\{\mathcal{S},\mathcal{A},r,P,\gamma\}$. Here, $\mathcal{S}$ and $\mathcal{A}$ denote the state space and action space, respectively; $r:(s,a)\rightarrow\mathbb{R}$ is the reward function; $P$ represents the state transition probability; and $\gamma$ is the reward discount factor. Specifically, the reward function is defined as $r(s,a)=\mathbb{E}[R_{t}\mid S_{t}=s,A_{t}=a]$. The policy is defined as $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$. Given the definition of the discounted return in reinforcement learning, denoted as
$U_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k},$ (7)
where $\gamma\in(0,1)$ is the discount factor and $R_{t}$ is the reward obtained at the step $t$ after the action $a_{t}$ of $s_{t}$. Let the state and action of time step $t$ be $s_{t}$ and $a_{t}$ respectively, the state-action value function (Q-function) at time step $t$ for a policy $\pi$ at state $s_{t}$ is given by
$Q_{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}\left[U_{t}|S_{t}=s_{t},A_{t}=a_{t}\right].$ (8)
And the state value function is defined as
$V_{\pi}(s_{t})=\mathbb{E}_{\pi}\left[U_{t}|S_{t}=s_{t}\right].$ (9)
Next, we introduce the performance difference lemma (PDL) [161]. Specifically, in PDL, $\mathbb{P}_{t}^{\pi}(s;s_{0})=\sum_{a}\mathbb{P}_{t}^{\pi}(s,a;s_{0})$ represents the probability that the policy $\pi$, starting from an initial state $s_{0}$, reaches state $s$ after $t$ time steps. Let $d^{\pi}_{s_{0}}(s,a):=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{t}^{\pi}(s,a;s_{0})$, and $d^{\pi}_{s_{0}}(s)=\sum_{a}d^{\pi}_{s_{0}}(s,a)$. Here, $d^{\pi}_{s_{0}}(s,a)$ represents the weighted frequency with which the state-action pair $(s,a)$ is visited in the trajectory when interacting with the environment under policy $\pi$. For the optimal policy $\pi^{*}$ and the predicted policy $\hat{\pi}$, PDL yields the following result
$|V_{\pi^{*}}(s_{0})-V_{\hat{\pi}}(s_{0})|=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi^{*}}_{s_{0}}}\left[\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[A_{\hat{\pi}}(s,a)\right]\right],$ (10)
where $A_{\hat{\pi}}(s,a)=Q_{\hat{\pi}}(s,a)-V_{\hat{\pi}}(s)$.
Expanding the innermost expectation on the right-hand side of $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[A_{\hat{\pi}}(s,a)\right]$, we obtain
$\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[A_{\hat{\pi}}(s,a)\right]=$ $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[Q_{\hat{\pi}}(s,a)\right]$ (11)
$-\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[V_{\hat{\pi}}(s)\right],$
where the second term is a constant independent of action $a$ so that $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[V_{\hat{\pi}}(s)\right]=V_{\hat{\pi}}(s)$. In fact, $V_{\hat{\pi}}(s)=\mathbb{E}_{a\sim\hat{\pi}(\cdot|s)}\left[Q_{\hat{\pi}}(s,a)\right]$. Based on this, we can obtain
$\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[A_{\hat{\pi}}(s,a)\right]=$ $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[Q_{\hat{\pi}}(s,a)\right]$ (12)
$-\mathbb{E}_{a\sim\hat{\pi}(\cdot|s)}\left[Q_{\hat{\pi}}(s,a)\right].$
@@
$|\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[Q_{\hat{\pi}}(s,a)\right]-\mathbb{E}_{a\sim\hat{\pi}(\cdot|s)}\left[Q_{\hat{\pi}}(s,a)\right]|\leq L\cdot\sqrt{\frac{\epsilon}{2}},$ (13)
where $L>0$ is the Lipschitz constant of the Q-function, $\epsilon=D_{\mathrm{KL}}(\pi^{*}(a|s)||\hat{\pi}(a|s))$. Combine Eq. 12 and Eq. 13, we get
$\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}\left[A_{\hat{\pi}}(s,a)\right]\leq L\cdot\sqrt{\frac{\epsilon}{2}}.$ (14)
Further, combining Eq. 10 and Eq. 14, we can get
$|V_{\pi^{*}}(s_{0})-V_{\hat{\pi}}$ $(s_{0})|\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\pi^{*}}}\left[L\cdot\sqrt{\frac{\epsilon}{2}}\right]$ (15)
$\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\pi^{*}}}\left[L\cdot\sqrt{\frac{D_{\mathrm{KL}}(\pi^{*}(a|s)||\hat{\pi}(a|s))}{2}}\right].$
The KL divergence term can be bounded using Eq. 3. In DRL, the target value function is externally defined, and a larger discount factor $\gamma$ implies that the objective places more emphasis on long-term outcomes. As $\gamma$ approaches 1, the gap between the predicted policy and the optimal policy tends to increase over an infinite time horizon. The Lipschitz constant $L$ characterizes the smoothness or complexity of the state-value function. Larger $L$ indicates a more complex objective, which in turn leads to a larger gap between the predicted and optimal policies.
In summary, Eq. 15 applies the existing KL divergence bound from conditional GDM distribution learning to derive a bound on the difference in state value functions between the learned and optimal policies. Unlike Eq. 6, which concerns one-shot solution generation, Eq. 15 considers externally defined reward objectives in MDP problems. Nevertheless, it similarly accounts for factors such as task complexity.
@@ V-B1 Strengths and Potential
As demonstrated by the above results, the learning errors of GDM in both one-shot optimization and MDP decision-making tasks admit bounded guarantees. This suggests that the distribution learning properties of GenAI enable convergence, and, when combined with existing theoretical frameworks [91, 110, 97, 98], help identify the various factors that contribute positively or negatively to the final learning outcomes. Consequently, the feasibility of applying GenAI to network optimization in these two paradigms is now supported by a solid theoretical foundation, which also offers guidance on practical design choices.
From an engineering perspective, GDM has demonstrated training stability in both paradigms and is making progress toward task generalization across diverse network optimization problems [93, 106]. Moreover, recent work has significantly reduced GDM’s computational and memory overhead [105], making it increasingly suitable for real-world deployment in various networking scenarios. In contrast, LPTMs show promising potential in tasks such as extracting user or operator intent from natural language problem descriptions and translating that into mathematical models for network optimization [163, 81], owing to their emerging capabilities in understanding and reasoning over mathematical content.
@@ V-B2 Flaws and Challenges
At a fundamental level, GenAI still faces inherent limitations such as output uncertainty, reliance on data-driven training, and autoregressive generation processes. When applied to network optimization, GDM’s solution and decision generation may degrade into regression modeling [164], offering limited generalization. LPTMs, meanwhile, internally compress and distinguish semantics in ways that diverge significantly from human cognition and often exhibit strong biases [165, 166, 167], as summarized in Table VII. The dominant autoregressive generation mechanism in LPTMs is also prone to error accumulation. While recent research increasingly explores reflective behaviors in these models to mitigate error accumulation [74], fundamental alternatives that challenge the autoregressive paradigm remain scarce [78].
@@
Empirically, a substantial body of newest research suggests that many of LPTMs’ capabilities in solving mathematical problems stem from pattern matching rather than genuine rule-based understanding [168, 46, 47, 48, 45]. That is, for tasks involving conceptual reasoning or arithmetic, LLMs often generate answers by recalling training samples similar in form to the given question, rather than by applying step-by-step logical rules. Moreover, the Transformer architecture has been shown to exhibit theoretical limitations in tasks involving function composition, i.e., decomposing and recombining subproblems or basic semantic units of the original problem [50, 51, 52, 53]. Similarly, GDMs rely on parallel sampling to generate multiple candidates and select higher-quality solutions or decisions. Like LPTMs, their performance is highly dependent on the completeness of training data, and they inherit GenAI’s broader difficulty in learning from long-tail distributions that can be particularly critical in the dynamic environments of network optimization.
TABLE VII: Recent critical work showing the limitations of GDMs and LPTMs.
{tblr}
colspec=X[0.05,l] X[0.42,c], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, References & Contributions
@@
[45] Demonstrates the performance limitations of state-of-the-art LPTMs in mathematical modeling within wireless communication tasks.
[46] Shows that current LLMs do not perform genuine logical reasoning on mathematical problems but instead attempt to replicate reasoning patterns observed during training.
[47] Confirms that reinforcement learning with verifiable rewards does not enhance the reasoning capabilities of LLMs, but only improves reasoning efficiency by reinforcing preferred solution paths.
[48] Reveals that existing LLMs fail to learn underlying mathematical principles and instead rely on memorizing patterns.
[50] Analyzes the limitations of the Transformer architecture in compositionality tasks that require multi-step reasoning, highlighting poor generalization to out-of-distribution inputs.
[51] Theoretically proves that a single Transformer layer cannot solve compositionality tasks with large feasible spaces, and that deeper Transformers may even struggle with basic arithmetic operations in such tasks.
[52] Provides a comprehensive analysis of the strengths and limitations of the representational capacity of the Transformer architecture.
[53] Theoretically compares the representational power of the encoder and decoder components in Transformers.
@@
[54] Empirically analyzes how dataset size, dimensionality, diversity, and neural hyperparameters affect memorization behaviors in GDMs.
@@
[55] Provides theoretical insights into how factors such as sample size, dimensionality, and training duration influence the generalization error of GDMs.
[158] Experimental results show that GDMs learn different distributions depending on the size of the training dataset.
[164] Evaluates the performance of conditional diffusion models on classification and regression tasks, indirectly suggesting a degeneration from generative to discriminative behavior in such tasks.
[165] Demonstrates that current LLMs compress semantic information in a way that differs significantly from humans, achieving information-theoretic compression at the cost of semantic detail.
[166] Experimental findings reveal conflicting internal representations of the same concept within LLMs.
[167] By analyzing intermediate representations, researchers uncover the internal reasoning flow of LLMs and identify intuition-driven outputs in mathematical question answering.

### v2

@@ V-A1 Bound of Solution Generation
Our goal is to relate the KL divergence between the GDM-generated solution distribution and the true target solution distribution to the expected objective value of the generated solutions. Due to the asymmetry of KL divergence, we adopt the forward KL divergence, rather than the reverse KL used in Eq. (2), to properly express the expected objective value under the generated distribution.
Specifically, we construct a target solution distribution induced by the objective function, denoted as $p_{0}(\mathbf{y}|\mathbf{x})\propto e^{-\alpha f(\mathbf{y},\mathbf{x})}$. Let $p_{0}(\mathbf{y}|\mathbf{x})=\frac{e^{-\alpha f(\mathbf{y},\mathbf{x})}}{Z(\mathbf{x})}$, where $Z(\mathbf{x})$ serves as the normalization constant or partition function, and $\alpha>0$ is a temperature parameter that controls the preference for optimal solutions. For large $\alpha$, the distribution concentrates around the optimal solutions; as $\alpha$ approaches zero, the distribution becomes nearly uniform. Then we assume the objective function $f$ is integrable with $\mathcal{X}$ and $\mathcal{Y}$ uniformly bounded. The normalization constant $Z$ can be expressed as $Z(\mathbf{x})=\sum_{\mathbf{y}\in\mathcal{Y}}e^{-\alpha f(\mathbf{y},\mathbf{x})}$ for discrete distributions and as $Z(\mathbf{x})=\int_{\mathcal{Y}}e^{-\alpha f(\mathbf{y},\mathbf{x})}d\mathbf{y}$ for continuous distributions, respectively.
@@
$=\alpha\mathbb{E}_{p_{\boldsymbol{\theta}}}\left[f(\mathbf{y},\mathbf{x})\right]\!+\!\log Z\!+\!\int_{\mathcal{Y}}p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})\log p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})d\mathbf{y}.$
With the entropy of $p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$ given as $\mathcal{H}(p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x}))=-\int_{\mathcal{Y}}p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})\log p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})d\mathbf{y}$, we can conclude that
@@
In Eq. (4), $\mathbf{y}^{*}$ denotes the optimal solution corresponding to the given $\mathbf{x}$. The normalization constant $Z$ depends only on the target solution distribution and $\mathbf{y}$, and is thus fixed under the given $\mathbf{x}$. The entropy term captures the uniformity of the model’s output distribution across solutions, higher entropy indicates a more uniform distribution, greater uncertainty, and a lower expected quality (i.e., higher objective value) of the generated solution. The parameter $\alpha$ characterizes both the distance between modes and the number of modes in the high-quality region of the target distribution. Intuitively, it reflects the complexity of the optimization problem. When $\alpha$ is small, the target distribution is more complex (i.e., more modes or closely spaced modes), implying higher problem complexity and a looser bound on the right side of Eq. (4). The KL divergence term is always non-negative. Although Eq. (4) is derived using the forward KL divergence, which differs from Eq. (2), the latter can still be interpreted as providing an upper bound for this term.
In summary, Eq. (4) quantifies the generalization boundary of the conditional GDM in one-shot optimization solution generation, and subsequently highlights several key factors that influence its generalization ability.
@@ V-A2 Bound of Policy Decision Generation
DRL trains a policy that takes actions based on the input state to receive rewards, aiming to maximize the expected cumulative discounted return. Our objective is to leverage the KL divergence bound between the optimal policy and the learned (predicted) policy, be described by Eq. (2), to quantify the gap between their corresponding value functions.
Consider an infinite horizon MDP $\mathcal{M}=\{\mathcal{S},\mathcal{A},r,P,\gamma\}$. Here, $\mathcal{S}$ and $\mathcal{A}$ denote the state space and action space, respectively; $r:(\mathbf{s},\mathbf{a})\rightarrow\mathbb{R}$ is the reward function; $P$ represents the state transition probability; and $\gamma$ is the reward discount factor. Specifically, the reward function is defined as $r(\mathbf{s},\mathbf{a})=\mathbb{E}[R_{t}\mid\mathbf{S}_{t}=\mathbf{s},\mathbf{A}_{t}=\mathbf{a}]$. The policy is defined as $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$. Accordingly, we denote the discounted return as $U_{t}$, the state–action value function (Q-function) as $Q_{\pi}(\mathbf{s}_{t},\mathbf{a}_{t})$, and the state value function as $V_{\pi}(\mathbf{s}_{t})$.
Next, we introduce the performance difference lemma (PDL) [159]. For the optimal policy $\pi^{*}$ and the predicted policy $\hat{\pi}$, PDL yields the following result:
$|V_{\pi^{*}}(\mathbf{s}_{0})-V_{\hat{\pi}}(\mathbf{s}_{0})|=\frac{1}{1-\gamma}\mathbb{E}_{\mathbf{s}\sim d^{\pi^{*}}_{s_{0}}}\left[\mathbb{E}_{\mathbf{a}\sim\pi^{*}(\cdot|\mathbf{s})}\left[A_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right]\right],$ (5)
where $A_{\hat{\pi}}(\mathbf{s},\mathbf{a})=Q_{\hat{\pi}}(\mathbf{s},\mathbf{a})-V_{\hat{\pi}}(\mathbf{s})$. Expanding the innermost expectation on the right-hand side of $\mathbb{E}_{\mathbf{a}\sim\pi^{*}(\cdot|\mathbf{s})}\left[A_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right]$, and transforming the expectation of the independent value into the form of a Q-function, we obtain
$\mathbb{E}_{\mathbf{a}\sim\pi^{*}(\cdot|\mathbf{s})}\left[A_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right]=$ $\mathbb{E}_{\mathbf{a}\sim\pi^{*}(\cdot|\mathbf{s})}\left[Q_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right]$ (6)
$-\mathbb{E}_{\mathbf{a}\sim\hat{\pi}(\cdot|\mathbf{s})}\left[Q_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right].$
@@
$|\mathbb{E}_{\mathbf{a}\sim\pi^{*}(\cdot|\mathbf{s})}\left[Q_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right]-\mathbb{E}_{\mathbf{a}\sim\hat{\pi}(\cdot|\mathbf{s})}\left[Q_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right]|\leq L\cdot\sqrt{\frac{\epsilon}{2}},$ (7)
where $L>0$ is the Lipschitz constant of the Q-function, $\epsilon=D_{\mathrm{KL}}(\pi^{*}(\mathbf{a}|\mathbf{s})||\hat{\pi}(\mathbf{a}|\mathbf{s}))$. Combine Eq. (6) and Eq. (7), we get
$\mathbb{E}_{\mathbf{a}\sim\pi^{*}(\cdot|\mathbf{s})}\left[A_{\hat{\pi}}(\mathbf{s},\mathbf{a})\right]\leq L\cdot\sqrt{\frac{\epsilon}{2}}.$ (8)
Further, combining Eq. (5) and Eq. (8), we can get
$|V_{\pi^{*}}(s_{0})-V_{\hat{\pi}}$ $(s_{0})|\leq\frac{1}{1-\gamma}\mathbb{E}_{\mathbf{s}\sim d_{s_{0}}^{\pi^{*}}}\left[L\cdot\sqrt{\frac{\epsilon}{2}}\right]$ (9)
$\leq\frac{1}{1-\gamma}\mathbb{E}_{\mathbf{s}\sim d_{s_{0}}^{\pi^{*}}}\left[L\cdot\sqrt{\frac{D_{\mathrm{KL}}(\pi^{*}(\mathbf{a}|\mathbf{s})||\hat{\pi}(\mathbf{a}|\mathbf{\mathbf{s}}))}{2}}\right].$
The KL divergence term can be bounded using Eq. (2). In DRL, the target value function is externally defined, and a larger discount factor $\gamma$ implies that the objective places more emphasis on long-term outcomes. As $\gamma$ approaches 1, the gap between the predicted policy and the optimal policy tends to increase over an infinite time horizon. The Lipschitz constant $L$ characterizes the smoothness or complexity of the state-value function. A larger $L$ indicates a more complex objective, which, in turn, leads to a larger gap between the predicted and optimal policies.
In summary, Eq. (9) applies the existing KL divergence bound from conditional GDM distribution learning to derive a bound on the difference in state value functions between the learned and optimal policies, and further identifies several key factors that influence this bound.
Remark (Generalization Bound).
When a GDM serves as a one-shot optimization solution generator or an MDP policy, its generalization error is bounded. This boundary is characterized by a KL-divergence-based bound on the generalization error of the GDM’s distribution-learning objective and highlights several influencing factors, including training duration, dataset size, problem complexity, and the discount factor.
@@ V-B1 Strengths and Potential
As demonstrated by the above results, the learning errors of GDMs in both one-shot optimization and MDP decision-making tasks admit bounded guarantees. This suggests that the distribution-learning properties of GenAI enable convergence and, when combined with existing theoretical frameworks [90, 109, 96, 97], help identify the various factors that contribute positively or negatively to final learning outcomes. Consequently, the feasibility of applying GenAI to network optimization across these two paradigms is now supported by a solid theoretical foundation that also offers guidance on practical design choices.
From an engineering perspective, GDMs have demonstrated training stability in both paradigms and is making progress toward task generalization across diverse network optimization problems [92, 105]. Moreover, recent work has significantly reduced GDM’s computational and memory overhead [104], making it increasingly suitable for real-world deployment in various networking scenarios. In contrast, LPTMs show promising potential for tasks such as extracting user or operator intent from natural-language problem descriptions and translating it into mathematical models for network optimization [161, 81], owing to their emerging capabilities for understanding and reasoning over mathematical content.
@@ V-B2 Flaws and Challenges
At a fundamental level, GenAI still faces inherent limitations, including output uncertainty, reliance on data-driven training, and autoregressive generation. When applied to network optimization, GDM’s solution and decision generation may degrade into regression [162], resulting in limited generalization. LPTMs, meanwhile, internally compress and distinguish semantics in ways that diverge significantly from human cognition and often exhibit strong biases [163], as summarized in Table IX. The dominant autoregressive generation mechanism in LPTMs is also prone to error accumulation. While recent research increasingly explores reflective behaviors in these models to mitigate error accumulation [74], fundamental alternatives that challenge the autoregressive paradigm remain scarce [78].
@@
Empirically, a substantial body of the newest research suggests that many of LPTMs’ capabilities in solving mathematical problems stem from pattern matching rather than genuine rule-based understanding [164, 46, 47, 48, 45]. That is, for tasks involving conceptual reasoning or arithmetic, LLMs often generate answers by recalling training samples similar in form to the given question, rather than by applying step-by-step logical rules. Moreover, the Transformer architecture has been shown to exhibit theoretical limitations in tasks involving function composition, i.e., decomposing and recombining subproblems or basic semantic units of the original problem [50, 51]. Similarly, GDMs rely on parallel sampling to generate multiple candidates and select higher-quality solutions or decisions. Like LPTMs, their performance is highly dependent on the completeness of training data, and they inherit GenAI’s broader difficulty in learning from long-tail distributions that can be particularly critical in the dynamic environments of network optimization.
TABLE IX: A set of recent works that shows the limitations or potential of GenAI and provides differing attitudes on the performance of generative optimization.
References Attitude Contributions Implications for Network Optimization
@@
[45] Negative Demonstrates the performance limitations of state-of-the-art LPTMs in mathematical modeling within wireless communication tasks. Highlights the need for task-specific adaptation and hybrid symbolic-neural optimization pipelines.
[46] Negative Shows that current LLMs replicate observed reasoning patterns instead of performing genuine logical reasoning. Suggests that semantic understanding and explicit constraint encoding are essential for reliable optimization.
[47] Negative Confirms that RLVR improves reasoning efficiency but not reasoning capability. Indicates that RLVR shows limited ability to improve GenAI’s inference for network optimization.
[48] Negative Reveals that existing LLMs rely on pattern memorization rather than learning mathematical principles. Implies the need for structure-aware optimization models that encode physical or mathematical priors.
[50] Negative Analyzes the failure of Transformers to generalize in compositional reasoning tasks. Suggests that compositional generalization remains a key bottleneck in scaled network optimization.
[51] Neutral Provides a comprehensive analysis of the representational strengths and weaknesses of the Transformer architecture. Provides reference for the upper and lower bounds of GenAI’s neural network backbone.
@@
[52] Neutral Empirically analyzes how dataset size, dimensionality, diversity, and hyperparameters affect memorization in GDMs. Emphasizes controlling data-model balance for stable generative optimization.
@@
[53] Neutral Provides theoretical insights into how sample size, dimensionality, and training duration influence GDM generalization error. Suggests generalization–efficiency trade-offs for GenAI in network optimization.
[76] Positive Enables theorem conjecturing and proving from axioms, demonstrating self-improvement and challenging theorem generation. Suggests self-reinforcing optimization mechanisms based on intrinsic reward signals.
[77] Positive Proposes a unified 5-element framework (scenario, parameters, variables, objectives, constraints) for optimizing generalization. Provides a generalizable paradigm for translating optimization problems into runnable codes.
[156] Neutral Shows that GDMs learn distinct distributions depending on training data size. Indicates the impact of sample quantity and diversity on GenAI learning results.
[163] Negative Demonstrates that LLMs achieve information-theoretic compression at the expense of semantic detail. Indicates that LPTM’s grasp of semantic details of encoded tokens is very limited compared to humans.
[165] Positive Theoretically proves that next-token prediction implicitly contains a universal reward model equivalent to one learned via inverse RL. Suggests leveraging intrinsic reward modeling for unsupervised network optimization.
[166] Positive Demonstrates that combining successful and failed experiences improves LLM reasoning via memory-aware test-time scaling. Encourages bidirectional experience replay for efficient online optimization.
[167] Positive Proposes Markovian Thinker, enabling stepwise RL reasoning under Markov assumptions, achieving significant efficiency gains. Suggests state-transition-aware network optimization strategies with bounded computational growth.